OpenStudy (anonymous):

simplify: y^2 -2y - 8 over y^2 - 7y +12 y-2/y-3 y-2/y+3 y+2/y+3 y+2/y-3

OpenStudy (anonymous):

y^2 -2y - 8 = (y - 4)(y + 2) y^2 - 7y +12 = (y - 3)(y - 4) y-4 cancel out